OpenStudy (anonymous):

Change the equation to quadratic form and solve it by the most efficient method. x − 8 = 20 over x x= (smaller value) x= (larger value)

OpenStudy (anonymous):

x^2-8x-20 = 0

OpenStudy (anonymous):

(x+10)(x-2)

OpenStudy (anonymous):

x = -10 or x = 2

OpenStudy (anonymous):

-10 < 2

OpenStudy (anonymous):

smaller value= -10? larger value=2? it's wrong

OpenStudy (anonymous):

x=-2 and x=10

OpenStudy (anonymous):

x^2-8x-20=0 x^2-10x+2x-20=0 x(x-10)+2(x-10)=0 (x+2)(x-10)=0 x=-2 or x=10

OpenStudy (anonymous):

thANK YOU!!!!!!!

OpenStudy (anonymous):

np